OpenStudy (anonymous):

take derivative of y= square root of (2x+1) by using limit definition?

OpenStudy (anonymous):

y'(x) = 1/sqrt(2 x+1)

OpenStudy (anonymous):

i know how to do derivative, but ca you show me the step for using limit definition?

OpenStudy (anonymous):

\[y \prime(x)=\lim_{h \rightarrow 0} (y(x+h)-y(x))/h)\]

OpenStudy (anonymous):

\[=\lim_{h \rightarrow 0} ((\sqrt{2(x+h)+1}-\sqrt{2x+1})/h)\] multiply by the conjugate \[=\lim_{h \rightarrow 0} ((\sqrt{2(x+h)+1}-\sqrt{2x+1})(\sqrt{2(x+h)+1}+\sqrt{2x+1})/h(\sqrt{2(x+h)+1}+\sqrt{2x+1})\] \[=\lim_{h \rightarrow 0} ((2(x+h)+1-2x+1)/h(\sqrt{2x+2h+1}+\sqrt{2x+1}))\] \[=\lim_{h \rightarrow 0} (2h/h(\sqrt{2x+2h+1}+\sqrt{2x+1}))\] \[=\lim_{h \rightarrow 0} (2/(\sqrt{2x+2h+1}+\sqrt{2x+1}))\] \[=2/(\sqrt{2x+2(0)+1}+\sqrt{2x+1})\] \[=2/(2\sqrt{2x+1})\] \[=1/\sqrt{2x+1}\]

OpenStudy (anonymous):

oops on the 3rd line i forgot brackets around the -2x+1 should be -(2x+1)

OpenStudy (anonymous):

thank you so much!